### v1

@@ 5.1 Maximum Singular Value Penalization
where $\lambda$ is the trade-off hyperparameter. The original Transformer (Vaswani et al., 2017) adopts multi-head self-attention to jointly attend to information from different subspaces at different positions. In line with that spirit, MSVP can also be added in a multi-head manner. Incorporating the summation over all the heads and layers, the overall training objective with multi-head MSVP is:
@@
where $\mathbf{W}_{l}^{*,h}$ denotes the $h^{th}$ head in the $l^{th}$ attention layer, and *** could be $Q,K$ or $V$. By constraining the maximum singular value of the weight matrices in the mapping, we can regulate the extent of output variations when subjected to attacks, thereby enhancing the model’s robustness. The computation of maximum singular values can be seamlessly integrated into the forward process.
In the next section, we show how to perform efficient computation of maximum singular values using the theoretically guaranteed (Thm. 5.1) power iteration algorithm.
@@
Algorithm 1 SpecFormer
@@
3:    Init $\bm{u}^{\ell}_{q/k/v},{\bm{v}}^{\ell}_{q/k/v}$
@@
6:A minibatch $\{(\mathbf{x}_{1},y_{1}),\ldots,(\mathbf{x}_{k},y_{k})\}$
@@
8:    Pass minibatch through the layers
@@
10:         ${\bm{v}}^{\ell}\leftarrow(W^{\ell})^{\top}\bm{u}^{\ell}$, $\bm{u}^{\ell}\leftarrow\mathbf{W}^{\ell}{\bm{v}}^{\ell}$,
11:         $\sigma^{\ell}\leftarrow\bm{u}^{\ell}\mathbf{W}^{\ell}{\bm{v}}^{\ell}$,
12:         Add $\lambda(\sigma^{\ell})^{2}$ to $\mathcal{L}_{msvp}$.
@@
15$\triangleright$ Backward:
16:Update $\mathbf{\theta}$ by backprop: $\mathcal{L}_{cls}+\mathcal{L}_{msvp}$.
Figure 1: SpecFormer with MSVP.
@@ 5.2 Power Iteration
Singular value decomposition (SVD) is the most direct way to calculate the maximum singular value. For any real matrix $\mathbf{A}\in\mathbb{R}^{m\times n}$, there exists a singular value decomposition of the form $\mathbf{A}=\mathbf{U}\mathbf{\Sigma}\mathbf{V}^{\top}$, where $\mathbf{U}$ is an $m\times m$ orthogonal matrix, $\mathbf{\Sigma}$ is an $m\times n$ diagonal matrix and $\mathbf{V}$ is an $n\times n$ orthogonal matrix. However, directly adopting SVD in MSVP will result in an additional time complexity of $\mathcal{O}(m^{2}n+n^{3})$, which poses a disastrous computational burden, particularly given the high hidden dimension of ViT and the vast number of attention layers used in the model.
In this section, we propose to adopt the power iteration algorithm as an alternative approach to efficiently calculate the maximum singular values for the $\mathbf{W}^{Q},\mathbf{W}^{K},\mathbf{W}^{V}$ weight matrices. The power iteration method (Burden et al., 2015) is a commonly used approach to approximate the maximum singular value with each iteration taking $\mathcal{O}(mn)$ time. By selecting an initial approximation vector $\bm{u}$ and ${\bm{v}}$, and then performing left and right matrix multiplication with the target matrix, we can obtain a reliable and accurate underestimate of the maximum singular value in a constant number of iterations. Moreover, as the algorithm is iterative in nature, we can incorporate the updates into the model’s update process, permitting us to efficiently estimate the maximum singular values with minimal cost. The following theorem provides a convergence guarantee of power iteration and the proof is in Appendix C. The complete training pipeline of SpecFormer is shown in Alg. 1.
Theorem 5.1 (Convergence guarantee of the power iteration method (Mises and Pollaczek-Geiringer, 1929)).
@@ 6.1 Setup
Datasets. We adopt four popular benchmark datasets: CIFAR10,CIFAR100 (Krizhevsky et al., 2009), ImageNet (Deng et al., 2009), and Imagenette (Howard, 2019). These datasets are commonly adopted in studies involving the robustness of vision models (Pang et al., 2022; Mo et al., 2022). CIFAR-10 and CIFAR-100 datasets each comprise $60,000$ images, categorized into $10$ classes and $100$ classes, respectively. ImageNet encompasses over $1.2$ million training images and $50,000$ test images, distributed across $1,000$ classes. Imagenette is a subset consisting of $10$ classes that are easy to classify, selected from the ImageNet. It is often employed as a suitable proxy (Mo et al., 2022; Wang et al., 2023) for evaluating the performance of models on the more extensive ImageNet.
Baselines. We compare SpecFormer with several notable baselines. Specifically, we evaluate SpecFormer against LipsFormer (Qi et al., 2023), which introduces Lipschitz continuity into ViT through a novel scaled cosine similarity attention (SCSA) mechanism and by replacing other unstable Transformer components with Lipschitz continuous equivalents. Additionally, we incorporate two other baseline models that implement Lipschitz continuity through different mechanisms: L2 multi-head attention (Kim et al., 2021) and pre-softmax Lipschitz normalization (Dasoulas et al., 2021). However, it is worth noting that the L2 multi-head attention approach requires $\mathbf{W}^{Q,h}=\mathbf{W}^{K,h}$, which is overly restrictive, potentially compromising the model’s representation power. We denote the L2 attention and Lipschitz normalization methods as L2Former and LNFormer, respectively.
Implementation details. In line with prior work (Mo et al., 2022), our evaluation spans across three distinct ViT backbones: the vanilla ViT (Dosovitskiy et al., 2020), DeiT (Touvron et al., 2021a), and ConViT (d’Ascoli et al., 2021). This approach enables us to perform a thorough validation and assess the general effectiveness of the MSVP algorithm. To ensure a fair comparison, we adopt the training strategy outlined in  (Mo et al., 2022). Specifically, we use the SGD optimizer with a weight decay of $0.0001$ and a learning rate of $0.1$ for $40$ epochs. By default, we apply both CutMix and Mixup data augmentation. We employ FGSM (Goodfellow et al., 2015) and PGD-2 (Madry et al., 2018) attacks for standard training, with an attack radius of $2/255$. For adversarial training, we use CW-20 (Carlini and Wagner, 2017) and PGD-20 (Madry et al., 2018) attacks, with an attack radius of $8/255$. In the case of the baseline models (Lipsformer, L2former, and LNFormer), we strictly adhere to the training protocols described in their respective papers.
@@ 6.2 Main Results
SpecFormer effectively controls the Lipschitz continuity of Transformers with minor modifications. From Table 1, 2 and 6, we can see that our proposed SpecFormer achieves the best results among all the competitors. Specifically, on standard training, our SpecFormer improves over the state-of-the-art robust ViT approaches by $\mathbf{2.79}$and $\mathbf{2.89}$in terms of robust accuracy on FGSM and PGD attacks, respectively, on average. Besides, we also improve the standard accuracy by $\mathbf{1.67}$%. Using adversarial training, our SpecFormer outperforms the best counterparts by $\mathbf{3.35}$and $\mathbf{2.58}$on CW and PGD attacks, respectively. Our method improves the clean accuracy of the ViT model by $\mathbf{3.80}$%. It significantly enhances the model’s robustness and its original performance.
We have more observations from the results. 1) Some baseline methods exhibit inferior performance when compared to vanilla ViTs, potentially attributable to the imposition of excessively stringent Lipschitz continuity constraints, which limit the model’s expressive capacity. 2) In contrast to other methods that ensure Lipschitz continuity in Transformers through self-attention mechanism modifications, our SpecFormer achieves this by adding a straightforward penalty term to the attention layers, without altering the original self-attention mechanism. SpecFormer’s simplicity and versatility allow seamless integration into various ViT architectures, preserving their flexibility. Experiment results on ImageNet and hyperparameter analysis is provided in Appendix F.1 and F.3.
@@
ViT-S LipsFormer (Qi et al., 2023) 71.13 31.48 4.17 40.05 9.92 1.36 86.80 36.60 30.20
L2Former (Kim et al., 2021) 79.65 39.98 13.39 53.20 15.35 5.92 92.80 58.00 37.80
LNFormer (Dasoulas et al., 2021) 75.82 33.72 7.75 48.81 13.04 7.27 92.00 49.00 41.80
ViT (Dosovitskiy et al., 2020) 87.09 45.56 22.35 63.52 19.82 7.01 94.20 72.60 48.00
@@
DeiT-Ti LipsFormer (Qi et al., 2023) 72.54 36.14 3.46 39.72 8.66 0.96 79.40 30.60 8.00
L2Former (Kim et al., 2021) 78.09 36.64 5.05 49.02 12.63 2.56 82.40 51.00 6.00
LNFormer (Dasoulas et al., 2021) 77.16 34.78 3.81 52.50 14.40 2.83 80.20 44.80 9.20
ViT (Dosovitskiy et al., 2020) 86.40 46.10 14.46 62.79 19.89 2.35 90.00 64.20 13.00
@@
ConViT-Ti LipsFormer (Qi et al., 2023) 79.71 38.47 7.09 48.08 10.84 1.57 90.60 44.40 24.60
L2Former (Kim et al., 2021) 81.33 40.17 12.76 49.86 13.86 2.03 93.40 62.20 30.80
LNFormer (Dasoulas et al., 2021) 75.48 28.67 3.56 51.13 11.62 2.41 84.20 36.20 15.00
ViT (Dosovitskiy et al., 2020) 87.78 48.89 20.26 64.68 22.94 4.96 93.20 68.80 40.80
@@
Standard CW-20 PGD-20 Standard CW-20 PGD-20 Standard CW-20 PGD-20
ViT-S LipsFormer (Qi et al., 2023) 41.54 24.20 27.50 24.08 10.47 13.04 50.00 31.00 34.80
L2Former (Kim et al., 2021) 63.22 33.55 35.78 37.20 13.69 15.65 84.80 54.60 54.60
LNFormer (Dasoulas et al., 2021) 56.04 29.92 32.74 26.37 9.84 11.63 81.00 46.80 48.80
ViT (Dosovitskiy et al., 2020) 71.76 34.34 35.49 36.45 11.97 12.89 89.60 62.80 62.40
SpecFormer (Ours) 72.73 31.84 31.90 41.46 12.80 13.54 91.60 67.00 67.00
DeiT-Ti LipsFormer (Qi et al., 2023) 39.72 24.05 26.77 23.09 9.82 12.03 39.00 27.60 28.40
L2Former (Kim et al., 2021) 60.85 34.29 36.63 36.67 14.23 16.48 77.40 41.40 43.20
LNFormer (Dasoulas et al., 2021) 54.32 29.85 32.96 28.65 11.36 13.94 72.20 39.00 42.20
ViT (Dosovitskiy et al., 2020) 71.71 37.15 38.74 40.89 15.25 17.40 79.00 40.60 41.40
SpecFormer (Ours) 80.03 48.52 51.10 44.48 16.36 18.21 82.20 46.20 46.00
ConViT-Ti LipsFormer (Qi et al., 2023) 56.83 32.27 35.04 31.50 14.56 17.17 60.80 34.00 38.20
L2Former (Kim et al., 2021) 39.36 23.84 26.42 16.53 8.06 9.65 10.00 10.00 10.00
LNFormer (Dasoulas et al., 2021) 49.03 28.68 31.68 29.12 13.00 15.65 72.20 39.00 42.20
ViT (Dosovitskiy et al., 2020) 53.09 30.87 33.63 31.54 14.65 17.24 68.00 41.40 43.60
SpecFormer (Ours) 67.05 38.72 41.58 37.92 14.50 16.78 86.60 47.20 46.20
@@
Figure 2: The analysis of MSVP. (a) Maximum singular value comparison between MSVP and the vanilla Transformer. (b)&(c) tSNE (Van der Maaten and Hinton, 2008) feature visualization.

### v2

@@ 5.1 Maximum Singular Value Penalization
where $\lambda$ is the trade-off hyperparameter. The original Transformer [53] adopts multi-head self-attention to jointly attend to information from different subspaces at different positions. In line with that spirit, MSVP can also be added in a multi-head manner. Incorporating the summation over all the heads and layers, the overall training objective with multi-head MSVP is:
@@
where $\mathbf{W}_{l}^{*,h}$ denotes the $h^{th}$ head in the $l^{th}$ attention layer, and * could be $Q,K$ or $V$. By constraining the maximum singular value of the weight matrices in the mapping, we can regulate the extent of output variations when subjected to attacks, thereby enhancing the model’s robustness. The computation of maximum singular values can be seamlessly integrated into the forward process.
In the next section, we show how to perform efficient computation of maximum singular values using a convergence-guaranteed power iteration algorithm.
@@
Algorithm 1 Detailed algorithm for SpecFormer
@@
3:     $\bm{u}^{\ell}_{q},{\bm{v}}^{\ell}_{q},\bm{u}^{\ell}_{k},{\bm{v}}^{\ell}_{k},\bm{u}^{\ell}_{v},{\bm{v}}^{\ell}_{v}\leftarrow$ initialize approximation vectors for $\mathbf{W}^{Q},\mathbf{W}^{K},\mathbf{W}^{V}$.
@@
6:Consider a minibatch $\{(\mathbf{x}_{1},y_{1}),\ldots,(\mathbf{x}_{k},y_{k})\}$ from training data.
@@
8:     Pass the minibatch through the layers, at the same time
@@
10:         for a sufficient number of times do $\triangleright$ One iteration is adequate
11:              ${\bm{v}}^{\ell}\leftarrow(W^{\ell})^{\top}\bm{u}^{\ell}$, $\bm{u}^{\ell}\leftarrow\mathbf{W}^{\ell}{\bm{v}}^{\ell}$, $\sigma^{\ell}\leftarrow(\bm{u}^{\ell})^{\top}\mathbf{W}^{\ell}{\bm{v}}^{\ell}$
12:              Add $\lambda(\sigma^{\ell})^{2}$ to maximum singular value loss $\mathcal{L}_{msvp}$.
@@
15:end for
16$\triangleright$ Backward:
17:Update parameters $\theta$ by backward propagation: $\mathcal{L}_{cls}+\mathcal{L}_{msvp}$.
@@ 5.2 Power Iteration
Singular value decomposition (SVD) is the most direct way to calculate the maximum singular value. For any real matrix $\mathbf{A}\in\mathbb{R}^{m\times n}$, there exists a singular value decomposition of the form $\mathbf{A}=\mathbf{U}\mathbf{\Sigma}\mathbf{V}^{\top}$, where $\mathbf{U}$ is an $m\times m$ orthogonal matrix, $\mathbf{\Sigma}$ is an $m\times n$ diagonal matrix and $\mathbf{V}$ is an $n\times n$ orthogonal matrix. However, directly using SVD in MSVP adds $\mathcal{O}(m^{2}n+n^{3})$ time complexity, creating a significant computational burden due to ViT’s high hidden dimensions and numerous attention layers.
In this section, we propose to adopt the power iteration algorithm as an alternative approach to efficiently calculate the maximum singular values for the $\mathbf{W}^{Q},\mathbf{W}^{K},\mathbf{W}^{V}$ weight matrices. The power iteration method [7] is a commonly used approach to approximate the maximum singular value with each iteration taking $\mathcal{O}(mn)$ time. By selecting an initial approximation vector $\bm{u}$ and ${\bm{v}}$, and then performing left and right matrix multiplication with the target matrix, we can obtain a reliable and accurate underestimate of the maximum singular value in a constant number of iterations. Moreover, as the algorithm is iterative in nature, we can incorporate the updates into the model’s update process, permitting us to efficiently estimate the maximum singular values with minimal cost. The convergence guarantee and proof for this algorithm can be found in Appendix C. The complete training pipeline of SpecFormer is shown in Alg. 1.
@@ 6.1 Setup
Datasets. We adopt four popular benchmark datasets: CIFAR10,CIFAR100 [27], ImageNet [16], and Imagenette [24]. These datasets are commonly adopted in studies involving the robustness of vision models [38, 34]. CIFAR-10 and CIFAR-100 datasets each comprise $60,000$ images, categorized into $10$ classes and $100$ classes, respectively. ImageNet encompasses over $1.2$ million training images and $50,000$ test images, distributed across $1,000$ classes. Imagenette is a subset consisting of $10$ classes that are easy to classify, selected from the ImageNet. It is often employed as a suitable proxy [34, 54] for evaluating the performance of models on the more extensive ImageNet.
Baselines. We compare SpecFormer with several notable baselines. Specifically, we evaluate SpecFormer against LipsFormer [42], which introduces Lipschitz continuity into ViT through a novel scaled cosine similarity attention (SCSA) mechanism and by replacing other unstable Transformer components with Lipschitz continuous equivalents. Additionally, we incorporate two other baseline models that implement Lipschitz continuity through different mechanisms: L2 multi-head attention [26] and pre-softmax Lipschitz normalization [13]. However, it is worth noting that the L2 multi-head attention approach requires $\mathbf{W}^{Q,h}=\mathbf{W}^{K,h}$, which is overly restrictive, potentially compromising the model’s representation power. We denote the L2 attention and Lipschitz normalization methods as L2Former and LNFormer, respectively.
Implementation details. In line with prior work [34], our evaluation spans across four distinct ViT backbones: the vanilla ViT [17], DeiT [50], ConViT [18] and Swin [29]. This approach enables us to perform a thorough validation and assess the general effectiveness of the MSVP algorithm. To ensure a fair comparison, we adopt the training strategy outlined in  [34]. Specifically, we use the SGD optimizer with a weight decay of $0.0001$ and a learning rate of $0.1$ for $40$ epochs. By default, we apply both CutMix and Mixup data augmentation. We employ FGSM [21] and PGD-2 [32] attacks for standard training, with an attack radius of $2/255$. For adversarial training, we use CW-20 [9] and PGD-20 [32] attacks, with an attack radius of $8/255$. We also evaluate performance using AutoAttack. In the case of the baseline models (Lipsformer, L2former, and LNFormer), we strictly adhere to the training protocols described in their respective papers.
@@ 6.2 Main Results
SpecFormer effectively enhances the adversarial robustness of ViTs with minor modifications. From Table 2, 4, 3, 5 and 6, we can see that our proposed SpecFormer achieves the best results among all the competitors. Specifically, under standard training, our SpecFormer improves over state-of-the-art robust ViT approaches by $\mathbf{8.96}$and $\mathbf{3.49}$in terms of robust accuracy on FGSM and PGD attacks, respectively, on average. Additionally, we enhance accuracy under AutoAttack for larger models by $\mathbf{2.36}$%. Furthermore, we improve standard accuracy by $\mathbf{2.20}$%.
Using adversarial training, our SpecFormer outperforms the best counterparts by $\mathbf{1.69}$and $\mathbf{1.26}$on CW and PGD attacks, respectively. For AutoAttack, we achieve an improvement of $\mathbf{0.50}$across all datasets and ViT variants, demonstrating stable robustness improvements. Our method also improves the clean accuracy of the ViT model by $\mathbf{3.06}$%, significantly enhancing both the model’s robustness and its original performance.
We have more observations from the results. 1) Some baseline methods exhibit inferior performance when compared to vanilla ViTs, potentially attributable to the imposition of excessively stringent Lipschitz continuity constraints, which limit the model’s expressive capacity. 2) AutoAttack is too powerful for small models under standard training, resulting in an accuracy of 00 in most cases. Therefore, we did not include the results for AA in Table 2. 3) In contrast to other methods that ensure Lipschitz continuity in ViTs through self-attention mechanism modifications, our SpecFormer achieves this by adding a straightforward penalty term to the attention layers, without altering the original self-attention mechanism. SpecFormer’s simplicity and versatility allow seamless integration into various ViT architectures, preserving their flexibility.
The results on ImageNet in Table 6 following the setup in [34] demonstrate that SpecFormer significantly outperforms its counterparts in both standard and robust accuracy using ViT-B as the backbone. Hyperparameter analysis is provided in Appendix E.1.
@@
ViT-S LipsFormer [42] 71.13 31.48 4.17 40.05 9.92 1.36 86.80 36.60 30.20
L2Former [26] 79.65 39.98 13.39 53.20 15.35 5.92 92.80 58.00 37.80
LNFormer [13] 75.82 33.72 7.75 48.81 13.04 7.27 92.00 49.00 41.80
TransFormer [17] 87.09 45.56 22.35 63.52 19.82 7.01 94.20 72.60 48.00
@@
DeiT-Ti LipsFormer [42] 72.54 36.14 3.46 39.72 8.66 0.96 79.40 30.60 8.00
L2Former [26] 78.09 36.64 5.05 49.02 12.63 2.56 82.40 51.00 6.00
LNFormer [13] 77.16 34.78 3.81 52.50 14.40 2.83 80.20 44.80 9.20
TransFormer [17] 86.40 46.10 14.46 62.79 19.89 2.35 90.00 64.20 13.00
@@
ConViT-Ti LipsFormer [42] 79.71 38.47 7.09 48.08 10.84 1.57 90.60 44.40 24.60
L2Former [26] 81.33 40.17 12.76 49.86 13.86 2.03 93.40 62.20 30.80
LNFormer [13] 75.48 28.67 3.56 51.13 11.62 2.41 84.20 36.20 15.00
TransFormer [17] 87.78 48.89 20.26 64.68 22.94 4.96 93.20 68.80 40.80
@@
Swin-Ti LipsFormer  [42] 84.77 0.05 35.94 60.60 0.68 14.06 89.20 12.40 30.60
L2Former [26] 97.86 44.33 54.69 88.24 12.76 29.69 99.40 66.00 95.20
LNFormer [13] 96.90 30.96 47.66 84.54 6.42 21.88 99.20 37.40 68.80
TransFormer [17] 97.51 42.91 53.91 87.89 14.16 25.00 99.60 61.40 95.31
SpecFormer (Ours) 98.25 58.15 53.64 89.07 16.24 28.74 99.40 73.00 96.80
Table 3: Results under standard training for larger models.
CIFAR10 CIFAR100 Imagenette
Standard PGD2 FGSM AutoAttack Standard PGD2 FGSM AutoAttack Standard PGD2 FGSM AutoAttack
LipsFormer [42] 83.32 13.83 40.52 8.26 59.05 7.97 18.11 4.00 90.80 29.80 32.00 1.20
L2Former [26] 90.67 18.80 47.01 8.33 75.03 11.71 24.87 5.83 97.60 40.40 77.40 5.20
ViT-B LNFormer [13] 86.68 14.20 40.67 5.99 65.90 8.35 17.77 4.09 92.00 31.00 38.80 3.20
TransFormer [17] 92.97 31.88 55.02 9.12 78.24 14.14 30.05 5.91 98.40 52.60 80.47 11.60
SpecFormer (Ours) 96.51 52.91 60.78 10.39 85.73 27.48 34.29 6.58 99.80 58.40 83.00 20.00
LipsFormer [42] 81.85 13.07 36.77 8.54 56.00 6.85 16.50 3.18 87.00 23.80 43.60 3.60
L2Former [26] 89.81 18.95 42.44 9.47 74.12 11.41 24.21 5.77 93.40 31.00 63.20 9.80
DeiT-S LNFormer [13] 86.48 13.78 39.32 6.94 67.01 8.40 20.05 4.17 91.40 31.80 54.20 8.40
TransFormer [17] 92.34 30.54 54.43 16.01 77.17 13.57 28.00 7.05 95.60 44.40 77.40 18.40
SpecFormer (Ours) 94.99 52.76 60.77 17.20 81.40 19.21 31.81 8.46 97.80 65.80 83.80 19.20
LipsFormer [42] 89.53 28.08 47.43 18.62 69.82 14.32 23.41 8.32 89.20 29.00 45.00 8.40
L2Former [26] 92.63 34.88 52.85 18.04 74.03 13.58 25.44 7.50 94.20 42.20 75.20 10.00
ConViT-S LNFormer [13] 87.84 17.96 42.68 9.54 67.79 8.84 19.29 4.43 89.00 16.20 36.80 6.60
TransFormer [17] 93.54 38.63 54.38 22.64 78.52 19.79 30.38 11.80 96.20 49.08 75.40 10.40
SpecFormer (Ours) 95.58 57.74 59.09 21.01 82.10 24.13 33.13 11.35 99.20 76.20 86.60 21.40
@@
Standard CW-20 PGD-20 AutoAttack Standard CW-20 PGD-20 AutoAttack Standard CW-20 PGD-20 AutoAttack
ViT-S LipsFormer [42] 41.54 24.20 27.50 23.74 24.08 10.47 13.04 9.71 50.00 31.00 34.80 37.00
L2Former [26] 63.22 33.55 35.78 36.35 37.20 13.69 15.65 18.78 84.80 54.60 54.60 62.20
LNFormer [13] 56.04 29.92 32.74 30.56 26.37 9.84 11.63 14.55 81.00 46.80 48.80 45.00
TransFormer [17] 71.76 34.34 35.49 46.32 36.45 11.97 12.89 24.29 89.60 62.80 62.40 60.60
SpecFormer (Ours) 72.73 31.84 31.90 47.07 41.46 12.80 13.54 23.25 91.60 67.00 67.00 64.60
DeiT-Ti LipsFormer [42] 39.72 24.05 26.77 23.07 23.09 9.82 12.03 9.54 39.00 27.60 28.40 23.80
L2Former [26] 60.85 34.29 36.63 34.03 36.67 14.23 16.48 17.24 77.40 41.40 43.20 48.20
LNFormer [13] 54.32 29.85 32.96 29.44 28.65 11.36 13.94 13.69 72.20 39.00 42.20 38.00
TransFormer [17] 71.71 37.15 38.74 43.60 40.89 15.25 17.40 20.89 79.00 40.60 41.40 56.40
SpecFormer (Ours) 80.03 48.52 51.10 45.36 44.48 16.36 18.21 23.49 82.20 46.20 46.00 58.20
ConViT-Ti LipsFormer [42] 56.83 32.27 35.04 30.75 31.50 14.56 17.17 14.15 60.80 34.00 38.20 33.40
L2Former [26] 39.36 23.84 26.42 22.21 16.53 8.06 9.65 9.03 10.00 10.00 10.00 10.00
LNFormer [13] 49.03 28.68 31.68 27.20 29.12 13.00 15.65 12.46 72.20 39.00 42.20 22.40
TransFormer [17] 53.09 30.87 33.63 37.78 31.54 14.65 17.24 14.47 68.00 41.40 43.60 39.20
SpecFormer (Ours) 67.05 38.72 41.58 35.30 37.92 14.50 16.78 18.23 86.60 47.20 46.20 36.20
Swin-Ti LipsFormer [42] 54.52 32.09 35.07 18.32 32.96 15.79 18.77 14.78 55.80 34.20 38.00 33.60
L2Former [26] 79.20 44.72 46.20 43.00 53.63 22.47 24.39 20.89 94.00 69.80 69.40 68.60
LNFormer [13] 77.39 43.33 45.13 41.68 52.90 22.48 24.70 20.76 93.00 67.60 68.00 66.80
TransFormer [17] 81.19 45.44 46.72 43.81 56.51 23.25 24.56 21.42 95.20 70.80 71.00 69.40
SpecFormer(Ours) 79.48 43.98 45.78 42.20 55.53 23.32 25.10 21.17 95.40 73.20 72.80 71.80
Table 5: Results under adversarial training for larger models.
Model Method CIFAR10 CIFAR100 Imagenette
Standard CW-20 PGD-20 AutoAttack Standard CW-20 PGD-20 AutoAttack Standard CW-20 PGD-20 AutoAttack
LipsFormer [42] 38.54 22.47 25.29 21.56 21.68 9.39 11.42 8.72 57.00 25.00 31.00 23.80
L2Former [26] 68.91 41.37 43.89 39.78 46.84 21.52 24.58 19.94 70.20 34.60 39.20 33.80
ViT-B LNFormer [13] 52.99 30.19 33.36 29.13 34.75 16.03 18.86 14.67 64.00 31.20 35.60 29.20
TransFormer [17] 82.70 50.17 52.65 48.27 61.18 28.33 30.16 26.14 91.40 65.80 67.40 64.40
SpecFormer (Ours) 87.22 51.32 52.55 48.80 54.91 24.66 27.73 22.80 89.80 63.20 64.80 62.00
LipsFormer [42] 41.90 25.00 27.79 24.03 25.17 11.30 13.54 10.40 47.60 31.00 34.40 30.80
L2Former [26] 67.86 39.80 42.73 38.13 45.38 21.12 23.87 19.66 84.60 55.00 55.60 53.60
DeiT-S LNFormer [13] 54.89 31.44 35.01 30.26 34.54 15.74 18.53 14.40 79.00 43.80 46.20 42.20
TransFormer [17] 81.37 49.22 51.83 47.34 58.73 27.15 29.47 25.13 91.40 66.60 66.20 65.60
SpecFormer (Ours) 83.61 50.03 52.01 47.97 61.76 27.84 29.70 25.48 93.40 67.40 67.00 65.80
LipsFormer [42] 38.99 23.59 26.40 22.72 15.92 7.81 9.54 7.39 59.40 33.80 38.40 32.60
L2Former [26] 32.50 23.15 24.42 22.58 23.61 10.82 13.29 10.22 10.00 10.00 10.00 10.00
ConViT-S LNFormer [13] 51.03 30.18 33.54 29.27 30.60 14.52 17.31 13.45 49.60 25.20 28.00 23.80
TransFormer [17] 63.84 36.63 39.75 35.21 29.02 13.21 15.73 12.34 89.20 62.80 63.40 60.20
SpecFormer (Ours) 67.71 38.50 41.49 36.85 32.95 15.26 17.94 14.16 92.40 67.20 66.80 65.40
Method Standard Training Adversarial Training
Standard FGSM PGD-2 Standard CW-20 PGD-20 PGD-100
LipsFormer [42] 65.76 20.87 3.77 45.04 18.91 21.13 20.83
L2Former [26] 77.40 37.12 6.89 51.24 24.85 27.04 26.95
LNFormer [13] 50.84 25.78 0.54 30.93 11.53 14.08 14.04
TransFormer [17] 79.11 41.45 10.79 60.81 30.92 32.58 32.35
SpecFormer (Ours) 80.04 43.51 11.59 62.30 31.87 32.82 32.56
@@
Figure 2: The analysis of MSVP. (a) Maximum singular value comparison between MSVP and the vanilla Transformer. (b)&(c) tSNE [31] feature visualization.